# Integrating Window-Based Correlated Decoding with Constant-Time Logical Gates for Large-Scale Quantum Computation

### v1

@@ Acknowledgements.
Supplemental Material
I correlated decoding within a window
In this section, we provide a explicit description of correlated decoding within a window. For CSS codes [6], the logical space is determined by stabilizer generators, denoted as $s_{x}^{(i)}$ and $s_{z}^{(i)}$, which represent $i$-th $X$-type and $Z$-type stabilizer generators, respectively. When running multiple rounds of syndrome extraction circuits, the error syndrome $\sigma_{n,k}(s_{x/z}^{(i)})\in\{+1,-1\}$ is defined as the measurement result of the stabilizer generator on the $n$-th logical qubit in the $k$-th round.
Next, the syndrome checks (or checks in brief) $C_{n,k}(s_{x/z}^{(i)})$, with value $c_{n,k}(s_{x/z}^{(i)})\in\{+1,-1\}$, are defined as the product of $\sigma_{n,k}(s_{x/z}^{(i)})$ and its previous result (if one exists), obtained by back-propagating the stabilizer operator through the circuit to the point where it was previously measured. Specifically, after the $k$-th round of syndrome extraction, if logical $I_{L}$, $X_{L}$, $Y_{L}$, or $Z_{L}$ gate is applied, then
$c_{n,k}(s_{x/z}^{(i)})=\sigma_{n,k}(s_{x/z}^{(i)})\sigma_{n,k-1}(s_{x/z}^{(i)}).$ (S1)
If logical $H_{L}$ gate is applied after the $k$-th round of syndrome extraction, then
$c_{n,k}(s_{x/z}^{(i)})=\sigma_{n,k}(s_{x/z}^{(i)})\sigma_{n,k-1}(s_{z/x}^{(i)}).$ (S2)
If logical CNOT gate between qubit $n_{1}$ and $n_{2}$ is applied after the $k$-th round of syndrome extraction, then
$c_{n_{1},k}(s_{x}^{(i)})$ $=\sigma_{n_{1},k}(s_{x}^{(i)})\sigma_{n_{1},k-1}(s_{x}^{(i)})\sigma_{n_{2},k-1}(s_{x}^{(i)}),$ (S3)
$c_{n_{2},k}(s_{x}^{(i)})$ $=\sigma_{n_{2},k}(s_{x}^{(i)})\sigma_{n_{2},k-1}(s_{x}^{(i)}),$
$c_{n_{1},k}(s_{z}^{(i)})$ $=\sigma_{n_{1},k}(s_{z}^{(i)})\sigma_{n_{1},k-1}(s_{z}^{(i)}),$
$c_{n_{2},k}(s_{z}^{(i)})$ $=\sigma_{n_{2},k}(s_{z}^{(i)})\sigma_{n_{2},k-1}(s_{z}^{(i)})\sigma_{n_{1},k-1}(s_{z}^{(i)}).$
In correlated decoding, each error event $E_{j}$ in the circuit is assigned an error probability $p_{j}$ and a set of checks $C_{j}=\{C_{n,k}(s_{x/z}^{(i)})|c_{n,k}(s_{x/z}^{(i)})=-1\text{ if }E_{j}\text{occurs}\}$. Each error event $E_{j}$ naturally corresponds to a hyperedge $e_{j}$ with a weight of $-\log p_{j}$, whose vertices correspond to the checks in $C_{j}$. Consequently, the correlated decoding problem is transformed into a minimum matching problem on a weighted hypergraph. According to Ref. [9], the solution to this problem can be exactly solved by a mixed-integer programming algorithm, known as the most-likely error (MLE) decoder, or approximately solved by hypergraph union-find (HUF) decoder [43, 44].
Next, let us describe how correlated decoding is executed within a window. First, we need to establish the relationship between each error event $E_{j}$ with its error rate $p_{j}$ and its corresponding hyperedge $e_{j}$. For each error event $E_{j}$, a noiseless circuit with the error event $E_{j}$ is simulated, and the set of checks with a value of -1, denoted as $C_{j}$, is recorded. However, in a window of limited size, some checks in $C_{j}$ may fall outside the window. Thus, in correlated decoding within a window, we remove checks that lie outside the current window and consider the remaining checks as the vertices of hyperedge $e_{j}$. If none of the vertices are within the current window, the error event $E_{j}$ is removed from the error event set of the current window. Additionally, hyperedges spanning both the commit and buffer regions will be removed after decoding. Therefore, for hyperedges spanning the commit and buffer regions of a window, we only consider the hyperedge $e_{j}$ and the corresponding error event $E_{j}$ in the current window, and remove $E_{j}$ from the error event set of the next window.
After each correlated decoding within a window, we preserve the hyperedges inside the commit region and those on the boundary of the commit and buffer region as reliable results. The syndrome information corresponding to these hyperedges is then eliminated, meaning the check values at the vertices of these hyperedges are flipped. The decoding of subsequent windows is then carried out, and the final output consists of all reliable hyperedges in each decoding as the complete result.
II size of temporal and spatial windows
For spatial windows, their size is immediately apparent. Each spatial window processes syndrome data from $\Theta(d)$ rounds on $n$ qubits at a time, so the size of correlated decoding remains acceptable. However, in TOQC circuits, a spatial window includes multiple $n$-qubit blocks. One might wonder whether the size of a spatial window remains acceptable. In the main text, we pointed out that the sizes of temporal and spatial windows are comparable. Here, we will more clearly demonstrate this conclusion.
To ensure fault tolerance, the distance between the commit region and the window boundary must be sufficiently large. This distance refers to the minimum number of hyperedges required to connect two vertices. In fact, this distance only depends on the size of the buffer region. For temporal and spatial windows, $n_{b}$ and $m_{b}$ can be set to $(d+1)/2$ and $(d+1)/2L$, respectively. This ensures that the probability of an error chain connecting the commit region to the window boundary is $\mathcal{O}(p^{(d+1)/2})$, which will not significantly contribute to the logical error rate $P_{L}$. On the other hand, the size of the commit region is not restricted. Typically, to output as many reliable decoding results as possible while keeping the overall window size manageable, the size of the commit region is set equal to that of the buffer region.
III Time-optimal quantum computation (TOQC) circuit
@@
Here, we explain additional details of the TOQC circuit mentioned in the main text. Generally, a TOQC circuit consists of two components, the teleportation circuit (Fig. S1) and the delayed choice circuit. When the input state $\psi$ is an $n$-qubit state, the operation $U_{i}$ is applied to a block composed of $n$ Bell pairs. Afterward, Bell basis measurements are performed, and based on the theory of teleportation, it can be easily verified that $U_{i}$ is successfully applied to the state $\psi$, up to a recovery operation $R$. When $U_{i}$ is a Clifford operation, the recovery operator $R$ is an element of the n-qubit Pauli group.
@@
In the Shor’s algorithm example mentioned in the main text, the non-Clifford gates in the TOQC circuit are typically Toffoli gates. The parallel execution of Toffoli gates can be achieved using delayed CZ gates with the similar methods in the TOQC, as discussed later.
Additionally, note that the TOQC circuit discussed in the main text is not entirely identical to the results in Ref. [20], which Ref. [20] can be regarded as a more general form of TOQC. In the general form, spatial window decoding is also applicable. This is because the general TOQC structure shares similarities with the one discussed in the main text. Specifically, the TOQC circuit can be divided into spatial blocks, and transversal CNOT gates between blocks only appear at the beginning or end of the circuit. As a result, the discussion in the main text is also valid for the TOQC described in Ref. [20].
Figure S1: Teleportation circuit in TOQC.
IV More details of numerical simulations
Figure S2: Details of the two examples in the numerical simulation. (a) Circuit and feed-forward sliding window for the first example. (b) Circuit for the second example. (c) Feed-forward spatial windows. (d) Parallel spatial windows.
Numerical simulations are conducted on two examples of sliding window applications in the main text. Here, we provide a clearer description of their details in Fig. S2. The logical qubits in the circuit are encoded on surface codes, enabling transversal CNOT gates and single-qubit Clifford gates from the set $\{X,Y,Z,H\}$. Fig. S2 a-b illustrate the specific circuits for both examples, where syndrome extraction is performed for single round after each logical gate. The division of windows is shown in Fig. S2 a and c-d.
In the second example, the circuit consists of five logical qubits and has a structure similar to TOQC with $L=(d+1)/2$. We tested the decoding process under two window architectures. In the feed-forward window architecture, each window contains syndrome data from two blocks. Correlated decoding is performed sequentially across four windows, with the window sliding by one block in the spatial direction each time. In the parallel window architecture, decoding is carried out in two steps: first, two disjoint spatial windows are decoded, followed by independent correlated decoding in the remaining buffer regions of the two parts.
We simulate the circuits under circuit-level noise model to obtained syndrome data. Specifically, the depolarizing Pauli noise channels are defined as follows:
$\mathcal{E}_{1}(\rho_{1})$ $=(1-p)\rho_{1}+(p/3)\sum_{P\in\{X,Y,Z\}}P\rho_{1}P,$ (S4)
@@ IV More details of numerical simulations
where $\rho_{1}$ and $\rho_{2}$ are single-qubit and two-qubit density matrices respectively. In the simulated circuits, we apply $\mathcal{E}_{1}$ after single-qubit gates, and idle operations, and $\mathcal{E}_{2}$ after two-qubit gates. Additionally, the measurement result and state initialization flips with the probability of $p$. We simulate the circuits in the Heisenberg representation [47], randomly inserting Pauli noise with corresponding probability. These Clifford circuits can be efficiently simulated in polynomial time, according to the Gottesman-Knill theorem [8, 48].
Then, correlated decoding is applied within each window. The decoding process uses an MLE decoder to obtain the exact solution for correlated decoding, implemented through the Gurobi solver [42]. One difference from the original correlated decoding literature is that we pre-merge several error events $\{E_{i}\}$ corresponding to the same hyperedge and assign an error probability based on the likelihood of an odd number of these events occurring. This approach accounts for the equivalence of error events. We expect the results after this pre-merging process to be closer to the exact solution for the decoding problem, i.e., most-likelihood decoding [49]. Additionally, the reduction in variables within the ${E_{i}}$ set implies a lower decoding time complexity.
@@
V implementation of ripple-carry adder and lookup table
In the main text, we discussed the two most time-consuming subroutines in Shor’s algorithm: the ripple-carry adder [38] and the lookup table (or QROM) [39]. Their specific circuit implementations can be found in Ref. [32] and Ref. [50]. The non-Clifford gate in the circuits of these two subroutines is the Toffoli gate. For surface codes and other quantum error correction codes with transversal CNOT gates, the Toffoli gate can be implemented using the gadget circuit in Fig. S3a, where the fixups are CZ and CNOT gates. To construct a TOQC circuit containing Toffoli gates, one can use the delayed CZ correction circuit shown in Fig. S3b, or equivalently implement a CCZ gate using the auto-CZZ ancilla state from Ref. [32].
@@
For the lookup table, another complexity lies in the implementation of multi-target CNOT gates. Without connectivity constraints, multi-target CNOT gates can be implemented using ancilla cat state (|$(\ket{0}^{\otimes n}+\ket{1}^{\otimes n})/\sqrt{2^{n}}$ followed by two layers of CNOT gates (only one layer on the computaional qubit), as shown in Fig. S4a. Additionally, the $n$-qubit cat state can be prepared in $\lceil\log n\rceil$ layers of CNOT gates, as shown in Fig. S4b. Note that in solid-state quantum systems, such as superconducting system, the time complexity of preparing an $n$-qubit cat state is $O(n)$, assuming only nearest-neighbor physical CNOT gates are allowed. Furthermore, multi-target CNOT gates cannot be heavily parallelized, as routing qubits will be occupied.
Figure S4: (a) Implementation of multi-target CNOT gate with ancilla cat state. (b) Circuit for preparing $n$-qubit cat state in $\lceil\log n\rceil$ layers.
VI A More Specific Estimate of the Space-Time Overhead for Shor’s Algorithm
@@
Our evaluation is not a complete reconstruction but rather focuses on comparing with the lattice surgery method, examining how fast transversal gates impact the overhead in various aspects. At the algorithm level, the analysis is the same as Ref. [21]. In other words, we maintain the same process flow of the algorithm as Ref. [21] but employ fast transversal gates rather than lattice surgery for the circuit implementation. In Ref. [21], several optimizations are made for the practical execution of Shor’s algorithm to factor a 2048-bit integer. Note that the parameters in this process are selected based on lattice surgery techniques for only superconducting systems. Thus, our analysis serves as a rough reference, leaving room for potential further optimization.
First, we clarify several important assumptions. It is assumed that ion trap systems allow arbitrary connectivity while maintaining the same error rate as superconducting systems. Besides, the time required for one round of QEC in ion trap systems is assumed to be $\gamma$ times longer than that of superconducting systems. Based on existing experimental data and theoretical projections, a $\gamma$ value between $10^{2}$ and $10^{3}$ is a reasonable assumption [26, 27, 28]. Moreover, the reaction time in ion trap systems is estimated to be $\eta$ times that of superconducting systems. In superconducting systems, the reaction time is mainly constrained by the time needed for classical processing and feedback, generally estimated to be around 10 $\mu$[32, 21]. For ion trap systems, the reaction time encompasses both the measurement and classical processing, and feedback, estimated to range between 20 and 200 $\mu$[35, 36, 37]. Therefore, $\eta$ is assumed to fall between 2 and 20. For simplicity, we assume that $\gamma$ and $\eta$ scale proportionally, i.e., $\gamma/\eta=50$. Additionally, the physical error rate is assumed to be $10^{-3}$, and the code distance for the computational qubits is set at $d=27$, which, as shown in Ref. [21], can provide a sufficiently low logical error rate.

### v2

@@ Acknowledgements.
Appendix A Time-optimal quantum computation (TOQC) circuit
@@
Here, we explain additional details of the TOQC circuit mentioned in the main text. Generally, a TOQC circuit consists of two components, the teleportation circuit (see Fig. 7) and the delayed choice circuit. When the input state $\psi$ is an $n$-qubit state, the operation $U_{i}$ is applied to a block composed of $n$ Bell pairs. Afterward, Bell basis measurements are performed Based on the theory of teleportation, it can be easily verified that $U_{i}$ is successfully applied to the state $\psi$, up to a recovery operation $R$. When $U_{i}$ is a Clifford operation, the recovery operator $R$ is an element of the $n$-qubit Pauli group.
@@
In the Shor’s algorithm example mentioned in the main text, the non-Clifford gates in the TOQC circuit are typically Toffoli gates. The parallel execution of Toffoli gates can be achieved using delayed CZ gates with the similar methods in the TOQC, as discussed in Appendix C.
Additionally, note that the TOQC circuit discussed in the main text is not entirely identical to the circuits in Ref. [24], which Ref. [24] can be regarded as a more general form of TOQC. In the general form, spatial window decoding is also applicable. This is because the general TOQC circuit shares similar structure with the one discussed in the main text. Specifically, the TOQC circuit can be divided into spatial blocks, and transversal CNOT gates between blocks only appear at the beginning or end of the circuit. As a result, the discussion in the main text is also valid for the TOQC described in Ref. [24].
Figure 7: Teleportation circuit in TOQC.
Appendix B Details of numerical simulations
Numerical simulations are conducted on two examples in the main text. We simulate the circuits under circuit-level noise model to obtained syndrome data. Specifically, the depolarizing Pauli noise channels are defined as follows:
$\mathcal{E}_{1}(\rho_{1})$ $=(1-p)\rho_{1}+(p/3)\sum_{P\in\{X,Y,Z\}}P\rho_{1}P,$ (5)
@@ Appendix B Details of numerical simulations
where $\rho_{1}$ and $\rho_{2}$ are single-qubit and two-qubit density matrices respectively. In the simulated circuits, we apply $\mathcal{E}_{1}$ after single-qubit gates, and idle operations, and $\mathcal{E}_{2}$ after two-qubit gates. Additionally, the measurement outcome and state initialization flips with the probability of $p$. We simulate the circuits in the Heisenberg representation [51], randomly inserting Pauli noise with corresponding probability. These Clifford circuits can be efficiently simulated in polynomial time, according to the Gottesman-Knill theorem [8, 52].
Then, correlated decoding is applied within each window. The decoding process uses an MLE decoder to obtain the exact solution for correlated decoding, implemented through the Gurobi solver [50]. One difference from the original correlated decoding literature is that we pre-merge several error events $\{E_{i}\}$ corresponding to the same hyperedge and assign an error probability based on the likelihood of an odd number of these events occurring. This approach accounts for the equivalence of error events. We expect the results after this pre-merging process to be closer to the exact solution for the decoding problem, i.e., most-likelihood decoding [53]. Additionally, the reduction in variables within the set $\{E_{i}\}$ implies a lower decoding time complexity.
@@
Appendix C Implementation of ripple-carry adder and lookup table
In the main text, we discussed the two most time-consuming subroutines in Shor’s algorithm: the ripple-carry adder [44] and the lookup table (or QROM) [45]. Their specific circuit implementations can be found in Ref. [38] and Ref. [54]. The non-Clifford gate in the circuits of these two subroutines is the Toffoli gate. For surface codes and other quantum error correction codes with transversal CNOT gates, the Toffoli gate can be implemented using the gadget circuit in Fig. 8a, where the fixups are CZ and CNOT gates. To construct a TOQC circuit containing Toffoli gates, one can use the delayed CZ correction circuit shown in Fig. 8b, or equivalently implement a CCZ gate using the auto-CZZ ancilla state from Ref. [38].
@@ Appendix C Implementation of ripple-carry adder and lookup table
For the lookup table, another complexity lies in the implementation of multi-target CNOT gates. Without connectivity constraints, multi-target CNOT gates can be implemented using ancilla cat state (|$(\ket{0}^{\otimes n}+\ket{1}^{\otimes n})/\sqrt{2^{n}}$ followed by two layers of CNOT gates (only one layer on the computaional qubit), as shown in Fig. 9a. Additionally, the $n$-qubit cat state can be prepared in $\lceil\log n\rceil$ layers of CNOT gates, as shown in Fig. 9b. Here we assume that the platform allows for all-to-all connectivity. Note that in solid-state quantum systems, such as superconducting system, the time complexity of preparing an $n$-qubit cat state is $O(n)$, assuming only nearest-neighbor physical CNOT gates are allowed. Furthermore, multi-target CNOT gates cannot be heavily parallel, as routing qubits will be occupied.
Figure 9: (a) Implementation of multi-target CNOT gate with ancilla cat state. (b) Circuit for preparing $n$-qubit cat state in $\lceil\log n\rceil$ layers.
Appendix D A More Specific Estimation of the Space-Time Overhead for Shor’s Algorithm
@@
Notably, our evaluation is not a complete reconstruction but rather focuses on comparing with the lattice surgery method, examining how constant-time transversal gates impact the overhead in various aspects. At the algorithm level, the analysis is the same as Ref. [25]. In other words, we maintain the same process flow of the algorithm as Ref. [25] but employ constant-time transversal gates rather than lattice surgery for the circuit implementation. In Ref. [25], several optimizations are made for the practical execution of Shor’s algorithm to factor a 2048-bit integer. Note that the parameters in this process are selected based on lattice surgery techniques for only superconducting systems. Thus, our analysis serves as a rough reference, leaving room for potential further optimization.
First, we clarify several important assumptions. It is assumed that ion trap systems allow arbitrary connectivity while maintaining the same error rate as superconducting systems. Besides, the time required for one round of QEC in ion trap systems is assumed to be $\gamma$ times longer than that of superconducting systems. Based on existing experimental data and theoretical projections, a $\gamma$ value between $10^{2}$ and $10^{3}$ is a reasonable assumption [29, 15, 30]. Moreover, the reaction time in ion trap systems is estimated to be $\eta$ times that of superconducting systems. In superconducting systems, the reaction time is mainly constrained by the time needed for classical processing and feedback, generally estimated to be around 10 $\mu$[38, 25]. For ion trap systems, the reaction time encompasses both the measurement and classical processing, and feedback, estimated to range between 20 and 200 $\mu$[41, 42, 43]. Therefore, $\eta$ is assumed to fall between 2 and 20. For simplicity, we assume that $\gamma$ and $\eta$ scale proportionally, i.e., $\gamma/\eta=50$. Additionally, the physical error rate is assumed to be $10^{-3}$, and the code distance for the computational qubits is set at $d=27$, which, as shown in Ref. [25], can provide a sufficiently low logical error rate for factoring 2048-bit integers.